This CD contains the freely distributable version of the ssh client for various Unixes and its associated file transfer program, scp. It also contains ssh clients for M$ Windows and Mac OS.

On the CD there are files README (for Unix users), README.TXT (For Windows users) and README.MAC (for MacOS users). These files have more details of the CD content and how to use it. On-line documentation may

be found at http://www-uxsup.csx.cam.ac.uk/CD/ .

On Solaris, Irix, Windows and MacOS a window showing the CD’s content should appear automatically on insertion of the CD.

Linux systems vary in their setup more widely than the commercial brands of Unix. On some Linux distributions there is an automounter running that will automatically mount the CD. On Red Hat Linux change directory to /misc/cd and see if the CD is mounted there. On Debian Linux, try /amnt/cdrom. Alternatively, users may be able to do the mount themselves. If the automounter does not seem to be mounting the CD and the directory /cdrom exists, try “mount /cdrom”. (Run “umount /cdrom” before ejecting the CD.)

If none of the above works you will have to manually mount the CD. For this you will need to contact the system’s administrator.

Queries by e-mail only, please, to: unix-support@ucs.cam.ac.uk

CD id: UCAM SSH CD 04